Monitoring Project - Introduction

The pendulum clock in Trinity College, Cambridge was installed in 1910. It is remarkably accurate, known to be better than one second per month. In 2009 a system was installed to monitor the "going". The work was done as part of a 4th year MEng project at the Engineering Department. A sensor on the pendulum detects the going of the clock, which is compared to the accurate time signal from a GPS receiver. For an entirely mechanical clock the accuracy is remarkable. Watch the clock strike twelve or hear it strike twelve at Singing from the Towers . Follow @clockkeeper on twitter. Today's and yesterday's weather in Cambridge, including wind. Here is a graph for the last 30 days of drift

A barometric compensator was installed at clock change - midnight 28 March 2010.
The clock is regulated as infrequently as possible so as to stay within +/- 5 seconds of the correct time. This enables the steady-state physics of the clock to be examined without unnecessary interference. The black dots • indicate that there is a comment relating to the time indicated. Blue arrows pointing down indicate that an adjustment was made to the Going of the clock by removing weights from the pendulum. Red arrows up indicate addition of weights to speed the clock up.

date [time GMT]Clock Diary
18 Mar 2018 [12:20:56]ADJUST: +700ms/day to 24000 in readiness for clock change next weekend  
18 Mar 2018 [12:05:06]winding by Cambridge Blue and Green Badge Guides: Deborah Vaudin, Fran De'Ath, Tim Pearce Higgins, Ruth Meyer and Helen Barrs  
16 Mar 2018 [16:17:00]Adjusted twist laser brightness. Not sure whether the rheostat had got knocked, or whether the moving of the prism had affected it, but the camera was underexposing, resulting in noisy measurements. (jpsa)  
16 Mar 2018 [12:51:49]visit by Nancy Hynes and Oliver Morton  
12 Mar 2018 [17:32:38]rotated the tilt meter by 180 degrees  
11 Mar 2018 [19:01:10]forgot to wind the clock earlier. Rap on knuckles. Wound and reset for 7pm.  
11 Mar 2018 [19:01:09]I got it all fixed in about 15 minutes: The clock had stopped at 5.44pm and it was restarted at 6.50pm so I had to advance the quarters by one notch. The Trinity Strike had stopped during the strike of two, so that needed advancing 5 notches. Fortunately the St John's strike was still in synch with the Trinity Strike. The clock had to be advanced by 1h6m, with all bells on silent. Then did a full wind  
09 Mar 2018 [12:43:17]ADJUST: +300ms/day to 23300 fine tuning  
09 Mar 2018 [12:35:36]advance by 80 seconds - forgot to do that last night  
09 Mar 2018 [00:20:00]ADJUST: +1000ms/day to 23000 after raising the laser prism up the pendulum shaft - this was to introduce up-down asymmetry in order to determine the sign of the twist measurement.  
04 Mar 2018 [00:37:40]ADJUST: -1400ms/day to 22000 fine tuning  
04 Mar 2018 [00:34:20]remove USB level (disturbed orange level)  
04 Mar 2018 [00:31:00]ADJUST: -5000ms/day to 23400 on account of removing the barocompensator  
04 Mar 2018 [00:26:00]release jack, observe further twist and tilt change  
04 Mar 2018 [00:21:00]jack up central beam to 200kg, observe further twist and tilt change  
04 Mar 2018 [00:15:00]jack up central beam by 100kg, observe twist and tilt change  
04 Mar 2018 [00:00:00]restart twist logger  
03 Mar 2018 [23:46:40]repeat twist experiment then restart pendulum  
03 Mar 2018 [23:45:00]installed new solid-state tilt sensor in parallel with the existing orange level  
03 Mar 2018 [23:07:30]stop pendulum, remove barometric compensator, replace twist prism, replace yellow-painted glass on laser screen with ground glass panel  
03 Mar 2018 [23:05:00]torsional disturbance to measure twist natural frequency while swinging and at rest  
03 Mar 2018 [23:00:00]arrived with Pip Mathews and John Aldridge for winding and to do some experiments. Removed heater from clock case. Put a curtain up in the window to block sunlight.  
03 Mar 2018 [17:00:00]put fan heater in clock case  
28 Feb 2018 [10:03:38]visit by Tom Besinger from Hainfeld and Frank Besinger from Berlin - just like old times!  
28 Feb 2018 [08:48:00]install new twist monitoring code, correct (at display time) erroneous pix->mrad scale factor used since 16-Jan (jpsa)  
26 Feb 2018 [09:30:00]with Stefan Savage, measuring some dimensions on the pendulum for some mods to be fitted on 24 March (clock change night). Switched off the level for a bit, and knocked the pendulum about accidentally.  
24 Feb 2018 [15:01:11]winding by Caitlyn Rayney and Charlie Vaughan - Perth and Melbourne respectively.  
17 Feb 2018 [12:15:00]replaced vellum screen on the laser twist sensor with a white polythene film. Thanks to John Aldridge for his help.  
17 Feb 2018 [10:35:00]noticed that two 200ms/day weights had fallen off the regulation platform. They were reinstated.  
17 Feb 2018 [10:07:45]Winding by Kelly Wing, Chloe Caron, Robert Spencer  
13 Feb 2018 [13:45:00]Oh no! Pigeon event. The clock stopped for 50 seconds. Is the pigeon wire not working?  
13 Feb 2018 [13:15:00]advance the clock by 50 seconds with Kristina Bifani  
09 Feb 2018 [22:28:48]winding with Rupert Thornely-Taylor and Emmanuel Dormy  
06 Feb 2018 [13:04:00]experiment, engaging the maintaining weight to see what effect it has on amplitude. Note that amplitude drops by about 0.09mrad.  
06 Feb 2018 [09:16:52]winding by Cameron Brick, David Brick, Mary Male, Kirsty Ross (and bump, again!), Adam Reeves Areeg Emarah  
02 Feb 2018 [14:30:01]ADJUST: -500ms/day to 28400  
29 Jan 2018 [21:25:00]winding by Roy Brander, Connie Brander and Smita Hope Gopisetty with Hilary Costello explaining the various temperature compensation, pressure compensation and general working of the clock - an enthusiastic and interested bunch  
29 Jan 2018 [14:41:07]ADJUST: +200ms/day to 28900 with Nick Breeze  
29 Jan 2018 [14:40:00]rotated the tilt sensor. It should now be giving readings arond +800mrads and the slope of plots vs tilt will be opposite sign  
29 Jan 2018 [14:38:20]shifted all 28900 to distance 200mm from pendulum centre  
27 Jan 2018 [14:02:00]shifted 23700 of the weights on the regulation to from 50mm to 150mm off-centre as an experiment to see if pendulum twist can be controlled.  
27 Jan 2018 [12:25:00]ADJUST: -5000ms/day to 28700 owing to the added mass of the regulation platform.  
27 Jan 2018 [12:05:00]clock restarted  
27 Jan 2018 [11:50:00]cavity probe reinstalled in cavity  
27 Jan 2018 [10:14:00]cavity probe removed and insulation spacers added. Laid down on the carpeted floor to check for zero offset. Note that by 11.45am the temperatures have settled down to the same value.  
27 Jan 2018 [09:45:00]pendulum stopped for access to pendulum case, plugging holes to get rid of draught. Also removed nut on Trinity Strike weight to give two dongs of extra time, hoping to get 7days 12hours between winds. Also fitted temporary platform to allow regulation weights to be placed up to 150mm off-centre as a means of regulating twist.  
27 Jan 2018 [09:41:50]09:41.50 close door to roof, End of experiment  
27 Jan 2018 [09:39:10]09:39:10 open door to roof,  
27 Jan 2018 [09:37:50]09:37.50 open clock case and remove shuttering around pendulum,  
27 Jan 2018 [09:36:10]09:36:10 open door to C5,  
27 Jan 2018 [09:34:30]09:34:30 close door to C5,  
27 Jan 2018 [09:32:15]09:32:15 pendulum case opened,  
27 Jan 2018 [09:29:15]09:29:15 hot water turned off,  
27 Jan 2018 [09:28:30]09:28:30 run hot water,  
27 Jan 2018 [09:26:50]09:26:50 close window,  
27 Jan 2018 [09:25:00]09:25:00 open window of C5,  
27 Jan 2018 [09:22:00]09:22:00 enter room C5 (Dr Hopkinson),  
27 Jan 2018 [09:16:00]Experiment to understand Bedmaker Effect: 09:16 enter clock room,  
26 Jan 2018 [17:35:00]end of experiment. [link] I reckon the shift of weight was about 0.25kg by 40mm - that's the same as 100kg by 0.1mm The pendulum bob is 100kg and the length is about 2m so the apparent tilt change is about 0.05mrad Look at x-y plot here of twist vs tilt (nb tilt volts is 1mrad/volt). [link] So it looks as if the 0.05mrad is in the right ballpark.  
26 Jan 2018 [15:00:00]experiment to see if shifting 20000ms worth of regulation weights makes any difference to twist. Looks promising  
25 Jan 2018 [21:28:11]winding by Robert Spencer, Khalid Aleen and Chloe Caron  
25 Jan 2018 [21:27:34]ADJUST: +1400ms/day to 33700  
24 Jan 2018 [10:46:00]improve chamber temperature logging: correct bottom-top order, log at hh:mm:00 times, log to full 1/16 C precision (jpsa)  
23 Jan 2018 [18:00:00]checked the pendulum chamber thermometers, cleaned out the muck at the bottom, accidentally knocked the pendulum. As a test, left pendulum case door open.  
18 Jan 2018 [09:14:28]winding by Chapel eventers, Eesha Khare, Frances Ding, Zhe Li, Cheryl Jaing, Peter Juhasz, Kirsty Ross (+Bump - big), Gui Fritas, Tobias Tan, Paul Johnson, Vijesh Bhute, Aga Machaj, Thomas Graf, Janko Ondras, Elizabeth Huang, Charlotte Keinhorst, Peggy Xu  
16 Jan 2018 [09:14:39]ADJUST: +200ms/day to 32300 with Chloe Caron  
15 Jan 2018 [18:56:36]ADJUST: -250ms/day to 32100 on account of lowering the prism slightly  
15 Jan 2018 [18:10:00]reinstalled the laser twist sensor and RPi. Much neater casing and wiring.  
11 Jan 2018 [12:11:40]ADJUST: -150ms/day to 32350  
11 Jan 2018 [12:10:00]winding with Harry Druiff and Len Howlett  
06 Jan 2018 [23:06:40]ADJUST: +300ms/day to 32500 fine tuning after raising the prism. Also refitting the laser spot light shield that had fallen off.  
06 Jan 2018 [23:05:00]winding with Eloise Hunt, Helen Hayward, Richard Hayward, Patrick Maxwell, Ed Maxwell, Wilke Grosche, Malte Grosche, Michael Watt, Alan Windle after the Twelfth Night Feast  
01 Jan 2018 [21:28:24]Debounced the edge detection for the pendulum & PPS sensors (jpsa)  
30 Dec 2017 [23:20:00]ADJUST: +1015ms/day to 32200 to account for the raising of the prism on the pendulum - and a top-up wind  
30 Dec 2017 [22:02:00]winding with Mark Irving and John Aldridge.  
30 Dec 2017 [20:32:00]Replaced twist laser unit with a sleeker version, and stopped pndulum to raise the prism into line with the new laser - does it work?  
23 Dec 2017 [13:05:00]winding with John Aldridge  
18 Dec 2017 [17:40:00]winding, and switch to Raspberry Pi with Rick Lupton and John Aldridge  
11 Dec 2017 [20:50:53]winding by Sander van der Linden, David Spiegelhalter, Anne Marthe van der Bles, Alex Freeman, Zsofia Szlamka, George Farmer, George Michael Pearson, Cameron Brick, Leila Finikarides  
10 Dec 2017 [19:06:07]clocklogger PC needed a "sudo reboot" and then "fsck /dev/sda1" (file system consistency check) which found many disk errors which were fixed.  
04 Dec 2017 [09:14:30]winding by Natalie Jones, Stephen Bayley, Annalise Higgins and Konrad Viebahn  
26 Nov 2017 [21:15:45]winding by Chloe Caron, William Baldwin, Ben Yass, Areeg Emarah  
19 Nov 2017 [10:13:40]winding by Chloe Caron, William Baldwin, Ben Yass  
11 Nov 2017 [23:30:42]ADJUST: +35ms/day to 31185 - by Chloe ... who is happy to take the blame.  
11 Nov 2017 [23:20:28]winding by Chloe Caron, Wang Yi, Chen Zhihang, Ben Yass, Wilfred Salmon, William Baldwin,  
04 Nov 2017 [12:41:34]ADJUST: -100ms/day to 31150  
04 Nov 2017 [12:39:07]winding by: William Baldwin, Rachel Onions, Phil Tower, Junhyuk Hwang, Philippa Nunn, Ben Yass, Max Vigdorchik, Haopeng Zhou, Chloe Caron,  
29 Oct 2017 [01:22:30]ADJUST: -50ms/day to 31250  
29 Oct 2017 [00:11:26]winding and time change by Ben Yass, Chloe Marie Leila Caron, Philippa Nunn, Rick Enns, Hilary Costello, John Aldridge, Chris Nickolaus, Alice McMichael, Andrew Gee  
23 Oct 2017 [20:22:56]winding by Carla Troyas Martinez, Sara Troyas Martinez, Giulio Crisanti, Adrian del Ser and Edward Karmazyn  
16 Oct 2017 [08:37:38]winding by Anisiya Parker, Ulyana Gumeniuk-Parker (past Fellow Commoner in Creative Arts) and Iain Parker  
08 Oct 2017 [21:25:54]winding by Eriks Esenvalds, Stephen 'sid' Layton, Paul Ashley - after evensong with Howells Cool. Reg  
01 Oct 2017 [11:28:25]winders from Christ's - (and a ring-in from Caius) Miriam Guth, Fin Brown, Emily Marr, Dom Yates, Vilda Markeviciute, Binbin Jin and Emily Morgan (the ring-in)  
01 Oct 2017 [11:28:01]ADJUST: +150ms/day to 31300  
23 Sep 2017 [23:09:53]winding after Erasmus Concert: Rose Bradshaw, Ellen Jameson, Dan Tidhar, Mary Earl, Michelle Rigozzi, Ian de Massini, Imogen Phillips, Chris White, Lucy Wilson, Edward Lilley, Johanna J. S. Finnemann, Patrick Welche, David Lavender, Alexander Blustin  
16 Sep 2017 [16:09:06]winding  
09 Sep 2017 [13:46:01]ADJUST: -100ms/day to 31150  
09 Sep 2017 [13:40:32]top-up wind by Richard Thomas, Joanna Pyman, Philippa Mann and Francesca Mann  
07 Sep 2017 [20:14:34]winding by Mark Burrell, Reece Oosterbeek and Laura Oosterbeek - a flock of Kiwis  
31 Aug 2017 [14:08:49]top-up wind by Jachin Chng Yee, marissa chng min, Chng Ying Elizabeth, Sarah Siah, Chng Kai Fong (matric '97) - visiting from Singapore  
30 Aug 2017 [12:05:12]top-up wind by Hilary Dennis, Oliver Dennis, Rosanne Hunt  
29 Aug 2017 [17:05:12]winding by membrers of the Trinity PostDoctoral Society  
22 Aug 2017 [17:05:12]winding by Philippa Matthews  
15 Aug 2017 [17:05:12]winding by Rick Lupton and Chris Nickolaus  
08 Aug 2017 [11:43:50]ADJUST: +200ms/day to 31250  
08 Aug 2017 [11:35:49]winding with TC PostDoc Society: Ruben Valbuena, Manon Kok, Sara Lopez-Gomollon,Jeongmin Choi, Francisco J Navarro, David Llewellyn-Jones, Gareth Richards and new Title-A Fellows George Roberts and Hannah Stern  
01 Aug 2017 [17:05:00]winding with Chris Nickolaus and friend  
25 Jul 2017 [17:05:00]winding by Rick Lupton and some of his research group  
20 Jul 2017 [09:00:00]visit with Kasia Ladds from Cambridge Enterprise, experiment to determine torsional period of pendulum: 10 cycles in 16.1sec so T = 1.61s  
19 Jul 2017 [12:09:02]ADJUST: -150ms/day to 31050 with John Aldridge helping to calibrate the tilt sensor  
18 Jul 2017 [18:34:02]winding with guests of the Cambridge Summer Music festival  
11 Jul 2017 [08:13:20]top-up wind, more experiments  
10 Jul 2017 [11:15:00]top-up wind, and experiments on twist vs angle  
08 Jul 2017 [22:42:27]winding by Jitka Stollova, Lynnea Shuck, Eleanor Cramer, Pip Matthews and Alice McMichael - after the Fairhaven Summer Concert  
01 Jul 2017 [11:15:00]top-up wind by Tony Howes, Catherine Howes, Oliver Howes, Max Howes amd Mimi - a flying vsit from Brisbane. Also Jonnie Hunt, Ellie Hunt and Alex Hunt  
30 Jun 2017 [17:13:14]winding by the committee of the Trinity Postdoctoral Society - Samantha Hajna, Simon Hoer, Barbara Herdy & Adina Feldman, President emerita of the Postdocs of Cambridge Society (PdOC)  
23 Jun 2017 [08:10:52]winding by Mechanics Lab technicians Tom Guilbride, Simon Marshall,Graham Smith, and Curator of the Whipple Museum of the History of Science Joshua Nall and also co-scouting parents Naomi Griffiths and Jonathan Spikey Griffiths  
15 Jun 2017 [20:13:05]wind for Rag: Hans Yu, Nicholas Jamieson, Maxim Kuvshinov, Jenna Foale, David Ren, Jose G. Aguilar, Petr Kungurtsev, Jocelino Rodrigues, Nguyen Anh Khoa Doan, Francesca De Domenico and Denis Busquet  
14 Jun 2017 [12:00:00]GCR=49.1s on lower layer,  
06 Jun 2017 [17:50:00]wind by Ryan Mulholland and Chris Nickolaus  
01 Jun 2017 [15:15:00]ADJUST: -150ms/day to 31200  
31 May 2017 [20:28:49]winding by Greg Winter, Andy Lim (matric '76), Hwee Hua Lim (matric G '78), Jane Corbett (matric '78), Lydia Heinrichs, Hans Yu  
24 May 2017 [17:43:13]Wenwen Li, Bingan Chen, Julia Clarke, Richard Clarke, David Boulton, Immanuel Kemp - winders from Arthur D Little, on a glorious sunny day  
22 May 2017 [21:03:20]ADJUST: -150ms/day to 31350 fine tuning  
22 May 2017 [15:46:40]rotating the prism - it wasn't quite straight  
22 May 2017 [15:25:00]ADJUST: -16800ms/day to 31500 on account of the added mass of the prism  
22 May 2017 [15:24:12]witnesses to fitting a prism and laser to the pendulum: Orlando Pope, James Pope, Margaret Pope, Eleanor Pope, Kenneth Pope, Anna Pope - believe it or not all related - and from AUCS Adelaide. Also Andrew Lantry from Worshipful Company of Clockmakers and Baba Soile MEng (almost!) We also refitted the temperature probe  
19 May 2017 [22:01:54]winding after a SPRNGG meeting with Renaud de Richter de Montpelier et Philip Davies de Birmingham!  
15 May 2017 [10:15:00]stopped pendulum to remove temperature probe from pendulum shaft - temperature sensors need to be cross-calibrated  
13 May 2017 [09:20:08]winding by Joe Sherwood - IIA Engineer, Lisa Hobley and Anthony Hobley - climate emergency compatriots!  
06 May 2017 [09:21:47]winding by Gael Vidal-Garner, Leila Vidal-Garner, Maia Vidal-Garner, Mandy Garner and IIA Engineers Ed Wheatcroft and James Pavelin  
29 Apr 2017 [15:59:35]winding by CSER delegates  
22 Apr 2017 [08:32:29]winding by 3C6 students Antoine Dupuis and George Hawkswell - first timers!  
21 Apr 2017 [20:15:00]What on earth happened here? [link]  
15 Apr 2017 [19:00:00]ADJUST: +200ms/day to 48300 and wind  
08 Apr 2017 [12:06:40]top-up wind  
08 Apr 2017 [12:05:00]ADJUST: +5000ms/day to 48100 to correct for extra weights, exactly as on 25 March  
08 Apr 2017 [12:00:00]I've reinstated the extra weights to the barocomp as they were on Sat March 25 - total eight extra weights - and I've superglued them in place so they don't fall off!  
07 Apr 2017 [18:01:00]winding with Catherine Aman and Tania Villares Balsa from Cambridge Enterprise - and some Spanish friends  
07 Apr 2017 [18:00:00]repositioning the fallen barometric compensator weight. No sign of mouse droppings...  
06 Apr 2017 [02:10:00]it looks as if another of the barometric compensator weights has fallen off. Will have to wait until winding tomorrow night to fix. I wonder if a mouse is the problem? Always seems to happen at 3am-ish.  
31 Mar 2017 [08:00:00]ADJUST: -300ms/day to 43100 and wind  
30 Mar 2017 [14:41:13]ADJUST: -4700ms/day to 43400 after one of the barometric compensator weights fell off! I removed the other one that was delicately poised. So the barocomp has a total of six extra weights (each 120g)  
29 Mar 2017 [09:43:20]For the record, here are instructions for Silencing the Clock [link] - useful for instance when the Choir is recording in the Chapel  
26 Mar 2017 [00:03:00]ADJUST: +200ms/day to 48100 fine tuning  
26 Mar 2017 [00:01:00]advance clock for summer time, done by Alice Wilcox  
25 Mar 2017 [23:51:00]winding by John Aldridge, Chris Nickolaus, Alice Wilcox, Graham Newman, Imogen Newman, Gary Walkow  
25 Mar 2017 [17:45:00]ADJUST: +7200ms/day to 47900 to correct for added weight to barocomp  
25 Mar 2017 [17:25:00]added 480g to barometric compensator (four 120g weights) to see if I can overcompensate it, as a check on its performance  
21 Mar 2017 [11:25:45]There was a power cut overnight. Restarted easily (no disk errors). Also a small top-up wind so we can get through to clock change at midnight on Saturday.  
18 Mar 2017 [11:30:00]winding with Graham Newman and Andy Burdon  
18 Mar 2017 [11:20:00]re-hanging the going tra 

Trinity College, Cambridge 2014